OpenStudy (anonymous):

The differential equation f'(x) = kf(x) has solutions of the form f(x) = Ce^(kx) (where k and C are constants) Produce an argument as to why this is the only solution. This should be an application of the constant Function Theorem.

OpenStudy (anonymous):

exponential are only function where derivatives are linear multiples of each other

OpenStudy (anonymous):

the constant function theorem is: if f'(x) = 0 for all x in (a,b), then f(x) = constant for all x in (a,b)

OpenStudy (anonymous):

how does that show that f(x) = Ce^(kx) are the only solutions? Is there a way to derive this from the theorem?

OpenStudy (anonymous):

also you can see when you solve it f'(x) ---- = k f(x) ln(f(x))= kx+ C exponentiate both side f(x)= e^kx+C= C e^kx

OpenStudy (anonymous):

is there a way I can come to this solution through the constant function theorem?

OpenStudy (anonymous):

I don't think constant function theorem address that

OpenStudy (anonymous):

it just say if f'(x)=0 then f(x)= some constant

OpenStudy (anonymous):

an example of an application of the theorem is: If f'(x) = g'(x) for all x in (a,b) then g(x) = f(x) + c for some constant c and all x in (a,b) let h(x) = g(x) - f(x), then h'(x) = g'(x) - f'(x) = 0 by the constant function theorem h(x) = g(x) - f(x) = c g(x) = f(x) + c

OpenStudy (anonymous):

is there something I can do that is similar to this?

OpenStudy (anonymous):

what's this constant function theorem, I can't find it anywhere

OpenStudy (anonymous):

the constant function theorem is: if f'(x) = 0 for all x in (a,b), then f(x) = constant for all x in (a,b)

OpenStudy (anonymous):

that's how I have it defined in my notes

OpenStudy (anonymous):

well I can't prove using your theorem , but whenever a function and their derivative set equal with constant multiples, the solution is always in exponential form

OpenStudy (anonymous):

how do you know that though?

OpenStudy (anonymous):

what's derivative e^x?

OpenStudy (anonymous):

what derivative of e^2x

OpenStudy (anonymous):

e^x and 2e^2x I think

OpenStudy (anonymous):

you see they are constant multiples of each other

OpenStudy (anonymous):

why wouldn't the C just be another k then?

OpenStudy (anonymous):

C and K are just constant

OpenStudy (anonymous):

but they should be the same constant shouldn't they?

OpenStudy (anonymous):

2 e^3x ----> 6 e^3x C=6 k=3

OpenStudy (anonymous):

oh alright I can see that now but, also even if e^x shows this property how do you know it is the only solution?

OpenStudy (anonymous):

no other function has that property

OpenStudy (anonymous):

but how can you show that this is the only one that works in this situation?

OpenStudy (anonymous):

well you gotta solve this differential equation y=c y'

OpenStudy (anonymous):

wouldn't you be solving f'(x) = kf(x)?

OpenStudy (anonymous):

ok y(x)=y'(x) k

OpenStudy (anonymous):

why is the k on the side with the derivative?

OpenStudy (anonymous):

doesn't have to be k is just a constant k y(x) =y'(x) y(x)= 1/k y'(x) y(x)=k y'(x) k is just constant , inverse of constant is still a constant

OpenStudy (anonymous):

ok I see how do I solve that equation then?

OpenStudy (anonymous):

I will write it as just y from now k y= y' k y = dy/dx ky dx = dy k dx= 1/y dy Integrate both sides kx+C= ln(y) exponentiate both sides e^(kx+C)= y e^C e^kx =y e^C is just as constant, we might as well right it as C C e^kx=y

OpenStudy (anonymous):

that does make sense now, thanks I'm still unsure of what my prof wanted me to do with the constant function theorem though. You don't have any more ideas about that do you?

OpenStudy (anonymous):

I look that up, didn't find much about it; but if f'(x)=0 ---> f(x)=C is true

OpenStudy (anonymous):

yeah I understand the theorem, I just can't figure out how this problem is an application of it